Quantitative Problem: Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis. Both projects' after-tax cash flows are shown on the time line below. Depreciation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firm's average project. Bellinger's WACC is 8%.

0 1 2 3 4
Project A -1,050 650 415 290 340
Project B -1,050 250 350 440 790

What is Project A's MIRR? Round your answer to two decimal places. Do not round your intermediate calculations.

%

What is Project B's MIRR? Round your answer to two decimal places. Do not round your intermediate calculations.

%
Answer #1

A:

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence future value of inflows=650*(1.08)^3+415*(1.08)^2+290*(1.08)^1+340

=$1956.0688

MIRR=[future value of inflows/Present value of outflows]^(1/time period)-1

=[1956.0688/1050]^(1/4)-1

=16.83%(Approx).

B:

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence future value of inflows=250*(1.08)^3+350*(1.08)^2+440*(1.08)^1+790

=$1988.368

MIRR=[future value of inflows/Present value of outflows]^(1/time period)-1

=[1988.368/1050]^(1/4)-1

=17.31%(Approx).